[figure: 9.7 cm]

QUANTUM TRANSFER-MATRIX SIMULATIONS OF THE SEMIMETALLIC S = 1/2 ANTIFERROMAGNETIC CHAIN Yb_{4}As_{3}

R. Matysiak^{a}, G. Kamieniarz^{b} J. Sznajd^{c} aInstitute of Fundamentals of Technology, University in Zielona Góra, Poland bInstitute of Physics, A. Mickiewicz University, Poznań, Poland

cInstitute of Low Temperature and Structure Research, Polish Academy of Sciences in Wrocław, Poland

Yb_{4}As_{3} is a new compound which represents the first example of an antiferromagnetic S = 1/2 spin chain where the soliton nonlinear excitation can be identified [1]. The experimental results give also a strong evidence for the existence of a staggered field alternating along Yb^{3+} chains induced by the Dzyaloshinski-Moriya interaction [2]. Applying the external magnetic field H perpendicular to the magnetic chain, the S = 1/2 anisotropic antiferromagnetic Heisenberg model with the Dzyaloshinski-Moriya interaction can be mapped on to the isotropic Heisenberg model with both uniform field H^{x} and staggered field H^{y}. In this communication we calculate specific heat and susceptibility of the model in question by the numerical quantum transfer-matrix method and compare the results with evailable experimental findings.

[1] M. Köppen et al., Phys. Rev. Lett. 82, 22 (1999).

[2] M. Kohgi et al., Phys. Rev. Lett. 86, 11 (2001).

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : R. Matysiak

Address for correspondence : Instytut Podstaw Techniki, Uniwersytet Zielonogórski, al. Wojska Polskiego 69, 65-762 Zielona Góra

Email address : matysiak@spin.amu.edu.pl